OpenStudy (anonymous):

please help !!!!!!! 7. What is the expression written using each base only once? (–4)^-6 · (–4)^7 4^13 –4 13 1

Parth (parthkohli):

Try using:\[a^m\cdot a^n = a^{m + n}\]

OpenStudy (anonymous):

it comes out to 1 so -4 because it cant have 1 as the base

Parth (parthkohli):

Yes :-)

OpenStudy (anonymous):

can you help with 1 more?

Parth (parthkohli):

Sure, sure :-)

OpenStudy (anonymous):

What is the value of y^(-5)/x^(-3) for x = 2 and y = –4?

OpenStudy (anonymous):

when i do it it comes out as -129/128 so would it be -1/128